10. A paper mill discharges chemicals into a river which pollutes the shores of a downstream resort area. The private

marginal cost, social marginal cost, and marginal benefit associated with the production of paper are given by

the following equations:

MCP = 5 +Q

MCS = 10 + 2Q

MB = 35-0.5Q

The marginal external cost at the allocatively efficient equilibrium is:

(a) 15

(b) 5

(c) 20

(d) 25

(e) 10
Answer #1

Answer 10

The allocatively efficient Quantity is the quantity where Marginal Social Benefit = Marginal Social cost.

Here there is only MB and thus MB = Marginal Social Benefit = 35 - 0.5Q

Marginal social cost = MCS = 10 + 2Q

Thus, Marginal Social Benefit = Marginal Social cost

=> 35 - 0.5Q = 10 + 2Q

=> Q = 10

Hence MCS = 10 + 2*10 = 30

Marginal private cost = MCP = 5 + 10 = 15

Marginal External Cost = Marginal Social cost(MCS) - Marginal Private cost(MCP)

= 30 - 15 = 15

Hence, the correct answer is (a) 15.